### v1

# ELMM: Efficient Lightweight Multimodal Large Language Models for Multimodal Knowledge Graph Completion

Abstract
Multimodal Knowledge Graphs (MKGs) extend traditional knowledge graphs by incorporating visual and textual modalities, enabling richer and more expressive entity representations. However, existing MKGs often suffer from incompleteness, which hinder their effectiveness in downstream tasks. Therefore, multimodal knowledge graph completion (MKGC) task is receiving increasing attention. While large language models (LLMs) have shown promise for knowledge graph completion (KGC), their application to the multimodal setting remains underexplored. Moreover, applying Multimodal Large Language Models (MLLMs) to the task of MKGC introduces significant challenges: (1) the large number of image tokens per entity leads to semantic noise and modality conflicts, and (2) the high computational cost of processing large token inputs. To address these issues, we propose Efficient Lightweight Multimodal Large Language Models (ELMM) for MKGC. ELMM proposes a Multi-view Visual Token Compressor (MVTC) based on multi-head attention mechanism, which adaptively compresses image tokens from both textual and visual views, thereby effectively reducing redundancy while retaining necessary information and avoiding modality conflicts. Additionally, we design an attention pruning strategy to remove redundant attention layers from MLLMs, thereby significantly reducing the inference cost. We further introduce a linear projection to compensate for the performance degradation caused by pruning. Extensive experiments on benchmark FB15k-237-IMG and WN18-IMG demonstrate that ELMM achieves state-of-the-art performance while substantially improving computational efficiency, establishing a new paradigm for multimodal knowledge graph completion.
ELMM: Efficient Lightweight Multimodal Large Language Models for Multimodal Knowledge Graph Completion
Wei Huang1, Peining Li1, Meiyu Liang1, Xu Hou1, Junping DU1, Yingxia Shao1, Guanhua Ye1, Wu Liu1, KangKang Lu1, Yang Yu1 1School of Computer Science, Beijing University of Posts and Telecommunications,
1 Introduction
Knowledge Graphs (KGs) represent real-world knowledge in a structured form using factual triples (head, relation, tail). In recent years, they have attracted considerable attention from both academia and industry, and have been widely applied to various downstream tasks such as information retrieval Yang (2020); Wang et al. (2024), event forecasting Huang et al. (2024); Gu and Krenn (2025), and recommendation systems Guo et al. (2020); Jiang et al. (2024). Multimodal Knowledge Graphs (MKGs) enhance traditional KGs by integrating information from multiple modalities, such as text and images, thereby enabling richer and more accurate knowledge representations. This multimodal fusion has shown to improve the performance of a wide range of intelligent systems Guo et al. (2024b). However, existing MKGs often suffer from incomplete information, which limits their effectiveness in practical applications. To address this challenge, the task of Multimodal Knowledge Graph Completion (MKGC) has emerged and has seen rapid development Gao et al. (2025); Zhang et al. (2024a); Chen et al. (2022); Shang et al. (2024). The core objective of MKGC is to leverage multimodal data to enhance the expressiveness of entity representations and uncover latent knowledge, thereby completing missing information in MKGs.
Large models, through extensive pre- and post-training on massive datasets, exhibit emergent capabilities and encode rich real-world knowledge Achiam et al. (2023); Guo et al. (2025a); Xiong et al. (2025), making them promising for knowledge graph completion (KGC). Recent studies have applied large language models (LLMs) to text-based KGC task. For example, KICGPT Wei et al. (2024) addresses the long-tail issue by re-ranking candidate triples from traditional methods, while MKGL Guo et al. (2024a) investigates whether LLMs can comprehend the structured representation of knowledge in the form of triples (head, relation, tail). However, current research largely overlooks multimodal inputs, leaving the potential of multimodal large language models(MLLMs) for MKGC underexplored.
In multimodal knowledge graphs (MKGs), entities are typically associated with various modalities of information, such as images and textual descriptions. For example, in datasets like FB15k-237-IMG Bollacker et al. (2008) and WN18-IMG Bordes et al. (2013), each entity is linked to 10 images along with corresponding textual descriptions, providing rich multimodal contextual information to complement the structured triples. State-of-the-art multimodal large language models (e.g., LLaVA-1.5Liu et al. (2024), Qwen2.5-VL Bai et al. (2025b), and MobileVLM Chu et al. (2023)) typically employ visual encoders, such as ViT Dosovitskiy et al. (2020), to encode different regions of an image into a sequence of image tokens. These image tokens are then concatenated with text tokens and jointly processed by MLLMs in a unified framework.
Figure 1: A description of each associated image region’s contribution to the entity Taylor Swift in a different link.
However, directly applying the aforementioned multimodal fusion paradigms to the task of multimodal knowledge graph completion presents several critical challenges. First, the number of image tokens per entity is large and often redundant. Standard visual encoders generate hundreds of tokens per image; with 10 images per entity, this results in thousands of tokens. Moreover, entities may exhibit different semantics under varying relational contexts, making the relevance of individual image tokens highly context-dependent. As shown in Figure 1, visual regions of the entity Taylor Swift vary in importance across different relations. Naively concatenating all image and text tokens introduces noise and modality conflicts, hindering semantic alignment. Second, the computational cost of processing such large token inputs is prohibitive. In multimodal large language models, the computational complexity of attention mechanisms typically grows quadratically with the number of tokens, while memory consumption increases at least linearly. This limits the model’s applicability in real-time and large-scale MKGC scenarios.
To address the aforementioned challenges, we propose ELMM, the first novel method to leverage Efficient Lightweight Multimodal Large Language Models (ELMM) for MKGC. First, we propose a Multi-view Visual Token Compressor (MVTC) based on multi-head attention mechanism to identify the most representative image tokens from both textual and visual views. This compressor adaptively compresses image tokens based on their semantic relevance to the textual description(containing entities and relations) and their intrinsic visual importance, compressing all image tokens from multiple input images into a fixed number of high-information tokens. This strategy significantly reduces visual noise and avoids modality conflicts, thereby enhancing both the inference efficiency and the representational capacity of the model. Second, through empirical observation, we identify a structural redundancy in the attention layers of MLLMs when applied to MKGC task—specifically, certain layers exhibit highly similar input and output representations, indicating limited contribution to overall performance. To mitigate this, we propose an attention pruning strategy that removes these redundant layers, effectively reducing computational overhead during inference. Additionally, we observe that the error introduced by the pruning strategy can be compensated by a linear projection, and in turn design a pipeline to initialize the linear projection.
Our contribution can be summarised as follows:
  • We propose Efficient Lightweight Multimodal Large Language Models (ELMM) for multimodal knowledge graph completion, a novel MKGC method that designs a Multi-view Visual Token Compressor (MVTC) based on multi-head attention mechanism to adaptively compress image tokens from both textual and visual views. This mechanism enables effective compression of multimodal inputs while mitigating modality conflicts and reducing noise. To the best of our knowledge, this is the first work to address the MKGC task using Multimodal Large Language Models.
  • We propose an attention pruning strategy that identifies and removes redundant attention layers in multimodal large language models (MLLMs) for the MKGC task, thereby reducing computational overhead. To mitigate pruning-induced errors, we apply linear projection as compensation.
  • Extensive experiments on the FB15k-237-IMG and WN18-IMG datasets demonstrate the superiority of ELMM in the MKGC task.
Figure 2: Overview of the ELMM architecture. From left to right: (1) ELMM processes input text and images containing incomplete triplet information; (2) MVTC adaptively compresses image tokens from both textual and visual views; (3) MLLM attention pruning and compensation mechanism.
2 Related Work
2.1 Multimodal Knowledge Graph Completion
Multimodal knowledge graph completion (MKGC) seeks to enhance the reasoning ability of knowledge graphs by jointly leveraging structural triples and multimodal entity information such as images and text. Early approaches Wang et al. (2019, 2021); Xie et al. (2016) typically used simple feature concatenation, but recent research has introduced more advanced techniques for fine-grained and adaptive multimodal fusion. Representative models include MKGformer Chen et al. (2022), which employs hybrid transformers with multi-level fusion; SGMPT Liang et al. (2024) introduces structural information based on MKGformer; LAFA Shang et al. (2024), which adopts link-aware fusion and neighbor aggregation; MyGO Zhang et al. (2024a), which tokenizes multimodal entity information for fine-grained representation; and NativE Zhang et al. (2024b), which addresses modality imbalance in real-world scenarios.
With the rapid advancement of multimodal large language models (MLLMs), these models have demonstrated substantial potential in task related to Multimodal Knowledge Graphs (MKGs). Although existing approaches have achieved notable progress in this area, most studies have yet to fully explore and leverage the capabilities of MLLMs. To address this gap, we propose ELMM, a novel method designed to harness the strengths of MLLMs.
2.2 Multimodal Large Language Models
With the rapid development of Multimodal Large Language Models(MLLMs), such as GPT-4 Achiam et al. (2023), LLaVA-1.5 Liu et al. (2024), Qwen2.5-VL Bai et al. (2025b), and Emu2 Sun et al. (2024), vision and language understanding have been deeply integrated within a unified semantic space. These models align visual and textual inputs through large-scale pretraining, enabling strong performance in visual question answering, captioning, and knowledge reasoning. Recent advances also focus on knowledge-enhanced pretraining Ma et al. (2024), multi-modal adapters Duan et al. (2024); Chen et al. (2025); Seputis et al. (2024), and efficient token selection to further improve performance in knowledge-intensive scenarios.
Despite the remarkable progress made by MLLMs across various multimodal tasks, they have yet to demonstrate significant breakthroughs in the task of MKGC. To address this limitation, we propose ELMM that incorporates several innovative mechanisms, including a Multi-view Visual Token Compressor (MVTC), an attention pruning strategy, and a multimodal knowledge reasoning completion layer. These components collectively enhance the capability of MLLMs in the MKGC task, leading to state-of-the-art performance in this domain for the first time.
3 Methodology
3.1 Preliminaries
3.1.1 Problem Formulation
A multimodal knowledge graph (MKG) is defined as a directed graph $MKG=(\mathcal{E},\mathcal{R},\mathcal{G},\mathcal{A}_{\mathcal{M}})$, where $\mathcal{E}$ and $\mathcal{R}$ denote the sets of entities and relations, respectively. The factual triples are represented as $\mathcal{G}=\{(e_{h},r_{h,t},e_{t})\mid e_{h},e_{t}\in\mathcal{E},r_{h,t}\in\mathcal{R}\}$. Each entity is associated with a set of multimodal attributes $\mathcal{A}_{\mathcal{M}}$, which includes textual $\mathcal{M}_{t}$ and visual $\mathcal{M}_{v}$ modalities. The task of multimodal knowledge graph completion (MKGC) aims to infer missing entities in $\mathcal{G}$. Given an incomplete triple $(h,r,?)$ or $(?,r,t)$, the goal is to predict the missing tail entity $t$ or $h$.
3.1.2 Multimodal Large Language Models
Generally, a multimodal large language model comprises the following components: a visual encoder and a textual tokenizer, which respectively segment images into multiple regions and encode it into a sequence of image token embeddings $I_{temp}\in\mathbb{R}^{NM\times E_{i}}$(NN represents the number of images, $M$ represents the region of each image, and $E_{i}$ represents the dimension), while simultaneously converting the input text into a series of text token embeddings $T_{t}\in\mathbb{R}^{K\times D}$($K$ is the number of text tokens and $D$ is the dimension). The image token embeddings are then projected into the same dimensional space as the text embeddings via a linear transformation, yielding $I_{t}\in\mathbb{R}^{NM\times D}$. These are concatenated with the text embeddings to form a unified multimodal token sequence $T\in\mathbb{R}^{(NM+K)\times D}$; A series of transformer blocks $\mathcal{M}$, which processes the multimodal token sequence and generates a corresponding sequence of hidden states for prediction;
$\mathbf{h}_{0:NM+K-1}^{m}=\mathcal{M}\left(t_{0:NM+K-1}\right)$ (1)
and a head layer, which maps each hidden state to a probability distribution $\mathbf{p}_{n}=head(\mathbf{h}_{NM+K}^{m})$.
3.2 ELMM Framework
The overall architecture of ELMM is depicted in Figure 2. Applying conventional multimodal large language models (MLLMs) directly to MKGC generates excessive image tokens (e.g., over 1,000 tokens from 10 entity-related images after encoding), introducing noise and exacerbating modality conflicts. To mitigate these issues, ELMM proposes a Multi-view Visual Token Compressor (MVTC) based on multi-head attention mechanism, which leverages multi-head attention to adaptively compress image tokens from both textual and visual views. MVTC effectively retains text-relevant information and key image details, thereby facilitating robust cross-modal alignment. Moreover, during MKGC, certain attention layers within the transformer blocks $\mathcal{M}$ exhibit high input-output similarity, indicating computational redundancy. ELMM addresses this by incorporating an attention pruning strategy that eliminates such redundant layers. A linear compensation module is then employed to mitigate approximation errors introduced by pruning, resulting in the pruned transformer blocks denoted as $\mathcal{M}^{\prime}$. Finally, to better support MKGC, ELMM replaces the traditional head layer with a multimodal knowledge reasoning completion layer, improving cross-modal fusion and generating a probability distribution over candidate entities.
3.3 Multi-view Visual Token Compressor
The overall workflow of Multi-view Visual Token Compressor (MVTC) is illustrated in the middle section of Figure 2. Considering that the importance of an image token can vary across different relational contexts (see Figure 1), MVTC first starts from the textual view and compresses image tokens using entity and relation tokens from text. This process retains the token representations that are most strongly associated with text information, thereby achieving modality alignment. Specifically, we concatenate the entity and relation token embeddings, denoted as $T_{\mathrm{entity}}$ and $T_{\mathrm{relation}}$, respectively, and apply max-pooling to obtain a fused textual representation $X_{t}\in\mathbb{R}^{1\times D}$. This representation $X_{t}$, together with the image token embeddings $I_{t}$, is used as input to a Multi-Head Attention (MHA) mechanism. The attention computation is defined as:
$Q_{i}=X_{t}W_{i}^{Q},K_{i}=I_{t}W_{i}^{K},Q_{i}\in\mathbb{R}^{1\times d},K_{i}\in\mathbb{R}^{MN\times d}$ (2)
$V=I_{t}W^{V},\quad V\in\mathbb{R}^{NM\times D}$ (3)
$\begin{split}Head_{i}&=Attention(Q_{i},K_{i},V)\\ &=softmax\bigg(\frac{Q_{i}K_{i}^{T}}{\sqrt{d}}\bigg)V,\quad Head_{i}\in\mathbb{R}^{1\times D}\end{split}$ (4)
$\begin{split}I_{text}=Concat(Head_{1},...,Head_{H}),I_{text}\in\mathbb{R}^{H\times D}\end{split}$ (5)
Here, $W_{i}^{Q},W_{i}^{K},W^{V}$ are trainable projection matrices; $d$ denotes the dimensionality of the queries and keys, where $d=D/H$, and $H$ is the number of attention heads. The obtained $I_{text}$ represents the result of compressing the image tokens from the textual view. Meanwhile, to ensure the effective preservation of core information in images, MVTC also performs compression on image tokens from the visual view. Specifically, to preserve the most representative global information from the images, we first extract the CLS tokens from multiple images and concatenate them to form:
$T_{\mathrm{CLS}}=\operatorname{concat}\left(\mathrm{CLS}_{1},\mathrm{CLS}_{2},\ldots,\mathrm{CLS}_{N}\right).$ (6)
We then apply a max-pooling operation over $T_{\mathrm{CLS}}$ to obtain a fused representation, denoted as $X_{i}$. Similar to the calculation process of $I_{text}$, we use $X_{i}$ and $I_{T}$ as the input of MHA to finally obtain $I_{image}$. The obtained $I_{image}$ contains the most important information of the image itself, representing the visual token information compressed from the visual view. We concatenate $I_{image}$, $I_{text}$, and $T_{t}$ to obtain $T$, and then reason through the pruned transformer blocks $\mathcal{M}^{\prime}$ to obtain hidden states $\mathbf{h}_{0:2H+K-1}^{m}$. As a result, the aggregated hidden states retain not only essential visual information but also visual cues semantically aligned with the textual modality.
3.4 Attention Pruning Strategy
Figure 3: Cosine similarity between the input and output of the attention layer in the training data of FB15k-237-IMG and WN18-IMG. The calculation method is to take 1,000 data points from the training set, perform inference, and calculate the average cosine similarity.
Due to the high computational cost of attention mechanisms, they often introduce significant inference latency. For instance, with an input sequence length of 512 tokens, the attention module alone can introduce up to 0.3437 seconds of latency per inference, accounting for approximately 77.9% of the total inference time (0.4410 seconds). To mitigate this overhead, we perform a redundancy analysis of the attention modules in MLLMs, such as LLaVA-1.5 Liu et al. (2024), during MKGC task. Specifically, we measure the cosine similarity between the input and output of each attention module to assess its redundancy. As illustrated in Figure 3, the upper layers of the model consistently exhibit high input-output similarity across various experimental settings. This observation suggests a significant degree of computational redundancy in certain attention module of MLLMs during MKGC task.
Based on this finding, we propose an Attention Pruning Strategy that selects the top-$K$ layers with the highest input-output cosine similarity and prunes their attention modules. However, the direct removal of attention mechanisms inevitably introduces a forward propagation error, denoted as $\varepsilon=x^{ori}-x^{pruning}$, where $x^{ori}$ is the original output and $x^{pruning}$ is the output after pruning. To compensate for the error we utilize a linear projection $x^{pruning}W_{c}$ to fit the error $\varepsilon$. Then, we propose a training-free initialization method for $W_{c}$. Furthermore, we provide the following theorem to guide the initialization of $W_{c}$:
Theorem 1.
The initialization of $W_{c}$ that minimizes the expected compensation error satisfies:
$W_{c}=V\Sigma^{+}U^{T}\mathbb{E}(\epsilon),$ (7)
where $U\Sigma V^{T}$ is the SVD decomposition of $\mathbb{E}(x^{pruning})$, $\Sigma^{+}$ denotes the pseudoinverse of $\Sigma$.
Proof.
Minimizing the expected compensation error corresponds to solving the following optimization problem:
$\min_{W_{c}}\|\mathbb{E}[x^{pruning}W_{c}-\varepsilon]\|^{2}.$ (8)
Since $W_{c}$ is independent of the data distribution, the expectation operator can be applied to $x^{pruning}$ and $\varepsilon$ separately. Thus, the problem reduces to:
$\min_{W_{c}}\left\|\mathbb{E}[x^{pruning}]W_{c}-\mathbb{E}[\varepsilon]\right\|^{2}.$ (9)
This is a standard least-squares problem with the solution:
$W_{c}=\mathbb{E}[x^{pruning}]^{+}\,\mathbb{E}[\varepsilon],$ (10)
where $\mathbb{E}[x^{pruning}]^{+}$ is the Moore–Penrose pseudoinverse of $\mathbb{E}[x^{pruning}]$. Given the SVD $\mathbb{E}[x^{pruning}]=U\Sigma V^{\top}$, the pseudoinverse is $\mathbb{E}[x^{pruning}]^{+}=V\Sigma^{+}U^{\top}$, leading to the final result:
$W_{c}=V\Sigma^{+}U^{\top}\,\mathbb{E}[\varepsilon].$ (11)
In practical applications, we randomly sampled 1,000 instances from the training dataset and employed the average values of $\bar{x}^{pruning}$ and $\bar{\epsilon}$ as estimators for $\mathbb{E}(x^{pruning})$ and $\mathbb{E}(\epsilon)$, respectively. Considering that errors in the early layers of the compensation model may propagate and affect the representations in subsequent layers, we adopted a bottom-up strategy to initialize $W_{c}$. Experimental results on the FB15k-237-IMG dataset demonstrate that, after applying linear compensation, the error of LLaVA-1.5 7B is significantly reduced from 276.5 (without compensation) to 4.7, corresponding to only 1.7% of the original error. This significant improvement underscores the effectiveness of our linear compensation method.
3.5 Multimodal Knowledge Reasoning Completion Layer
To fit the multimodal knowledge graph completion task, we employ a multimodal knowledge reasoning completion layer to compute the probability distribution over candidate entities. The process is defined as follows:
$E_{m}=[E_{image},E_{entity},E_{relation}],S\in\mathbb{R}^{1\times 3D}$ (12)
$E_{m}^{\prime}=LinearLayer(E_{m}),E_{m}^{\prime}\in\mathbb{R}^{1\times 3D}$ (13)
$p=LinearLayer(E_{m}^{\prime}),p\in\mathbb{R}^{1\times r}$ (14)
Here, $E_{image}$ is obtained by max-pooling over the hidden states of image modality hidden states. $E_{entity}$ and $E_{relation}$ correspond to the hidden states of the entity and relation tokens. These are concatenated to form the multimodal representation $E_{m}$. The fused representation is first processed by a linear layer and then projected to obtain the probability distribution $p\in\mathbb{R}^{1\times r}$ of the candidate entities, where $\mathit{r}$ represents the number of entities.
3.6 Optimization
To optimize ELMM, we employ a contrastive learning objective, defined as follows:
$\begin{split}\mathcal{L}&=\sum_{(e_{i},r_{k},e_{j})\in\mathcal{T}_{train}}\bigg[-\log(p_{e_{j}})\\ &\quad+\frac{1}{|\mathcal{N}_{neg}(e_{j})|}\sum_{e_{neg}\in\mathcal{N}_{neg}(e_{j})}\log(1-p_{e_{neg}})\bigg].\end{split}$ (15)
Dataset #Rel. #Ent. #Train #Dev #Test
FB15k-237-IMG 237 14,541 272,115 17,535 20,466
WN18-IMG 18 40,943 141,442 5,000 5,000
Table 1: Statistics of Dataset.
Here, $\mathcal{N}_{neg}\left(e_{j}\right)=\left\{e_{neg}\cup e_{i}\mid e_{neg}\neq e_{j},e_{neg}\in\mathcal{E}\right\}$ denotes the set of negative samples for the target entity $e_{j}$. Since the prediction of tail entities in the form of $(e_{i},r_{k},\ ?)$ often results in a bias where the head entity $e_{i}$ is assigned disproportionately high confidence, we incorporate a self-denoising strategy by treating the head entity $e_{i}$ as a hard negative sample. The loss function $\mathcal{L}$ is in a form of a binary cross-entropy.
4 Experiments
Model Name FB15k-237-IMG WN18-IMG
MR Hits@1 Hits@3 Hits@10 MR Hits@1 Hits@3 Hits@10
Multi-modal KGC Models
IKRL(IJCAI 2017) 298 19.4 28.4 45.8 596 12.7 79.6 92.8
TransAE(IJCNN 2019) 431 19.9 31.7 46.3 352 32.3 83.5 93.4
RSME(ACMMM 2021) 417 24.2 34.4 46.7 223 94.3 95.1 95.7
KG-BERT(2019) 153 - - 42.0 58 11.7 68.9 92.6
VisualBERT(2019) 592 21.7 32.4 43.9 122 17.9 43.7 65.4
VILBERT(ICLR 2020) 483 23.3 33.5 45.7 131 22.3 55.2 76.1
MKGformer(SIGIR 2022) 221 25.6 36.7 50.4 28 94.4 96.1 97.2
LAFA(AAAI 2024) 136 26.9 39.8 55.1 25 94.7 96.5 97.7
NativE(SIGIR 2024) 149 25.4 38.6 54.2 35 94.2 95.8 97.1
SGMPT(ACMMM 2024) 238 25.2 37.0 51.0 29 94.3 96.6 97.8
MyGO(AAAI 2025) - 19.0 28.9 44.7 70.6 93.7 94.1
MPIKGC(COLING 2024) - 24.4 35.8 50.3 - - -
AdaMF-MAT(COLING 2024) 23.1 35.0 49.1 73.6 94.3 95.8
Uni-modal KGC Models
KICGPT (EMNLP 2023) 154 32.7 44.8 55.4 25 93.5 95.6 96.8
MKGL (NeurIPS 2024) 172 32.5 45.4 59.1 31 93.6 95.4 96.3
K-ON (AAAI 2025) 144 33.2 45.7 59.8 26 94.3 95.7 97.2
ELMM(Ours) 112 34.1 47.3 60.7 11 95.7 97.5 98.7
{*} indicates that the result comes from SGMPT Liang et al. (2024).
Table 2: Models performance on the FB15k-237-IMG and WN18-IMG datasets. Bold and underline text indicate the best and second-best performance, respectively.
4.1 Datasets
We evaluate the performance of ELMM on two publicly available multimodal knowledge graph datasets: FB15k-237-IMG Bollacker et al. (2008) and WN18-IMG Bordes et al. (2013). Each dataset comprises three modalities: (1) structured knowledge graph triples, (2) textual descriptions of entities, and (3) a set of 10 associated images per entity. Detailed statistics for both datasets are presented in Table 1.
4.2 Setting
We adopt LLaVA-1.5 7B Liu et al. (2024) as the base multimodal large language model (MLLM), incorporating Low-Rank Adaptation (LoRA) Hu et al. (2022) techniques into both the query and value layers to enable parameter-efficient fine-tuning. All experiments are conducted using eight NVIDIA A100 GPUs, and both training and inference are implemented with the PyTorch framework. The parameter $K$ in the attention pruning strategy is always set to 16, indicating that 16 attention modules in the MLLM are pruned and replaced with linear projections.
To evaluate the effectiveness of ELMM on the Multimodal Knowledge Graph Completion (MKGC), we adopt standard evaluation metrics, including Hits@k (including Hits@1, Hits@3, and Hits@10) to assess ranking precision at different thresholds, and Mean Rank (MR) to measure the average ranking position of the correct target entities among the predictions.
For a comprehensive performance comparison, we benchmark our model against two representative categories of methods: (i) multimodal knowledge graph completion methods, and (ii) uni-modal knowledge graph completion methods based on large language models (LLMs). The former category includes IKRL Xie et al. (2016), TransAE Wang et al. (2019), RSME Wang et al. (2021), KG-BERT Yao et al. (2019), VisualBERT Li et al. (2019), ViLBERT Su et al. (2019), MKGformer Chen et al. (2022), LAFA Shang et al. (2024), NativE Zhang et al. (2024b), SGMPT Liang et al. (2024), MyGO Zhang et al. (2024a), MPIKGC Xu et al. (2024), and AdaMF-MAT Zhang et al. (2024c). The LLM-based uni-modal knowledge graph completion baselines include KICGPT Wei et al. (2024), K-ON Guo et al. (2025b) and MKGL Guo et al. (2024a). the traditional most powerful multimodal methods
4.3 Comparison with State-of-the-art
We conduct a comprehensive comparison between the proposed ELMM method and sixteen state-of-the-art models across two benchmark datasets. As shown in Table 2, ELMM consistently outperforms all competing methods across all evaluation metrics. Notably, on the FB15k-237-IMG dataset, ELMM demonstrates substantial improvements over most powerful multimodal approaches, achieving relative gains of 26.8%, 18.8%, and 10.2% in Hits@1, Hits@3, and Hits@10, respectively. These results strongly validate the powerful potential of multimodal large language models (MLLMs) in the task of multimodal knowledge graph completion. Furthermore, compared to uni-modal methods based on large language models, ELMM exhibits clear advantages on both datasets, underscoring its effectiveness and robustness in extracting and integrating multimodal information.
4.4 Ablation Study
To evaluate the effectiveness of each component in ELMM, we conduct a series of ablation studies by removing or modifying key modules of the model. Specifically: w/o Image removes the visual views image token representations ($I_{image}$); w/o Text removes the textual view image token representations ($I_{text}$); w/o MVTC removes MVTC and does not compress image tokens; w/o pruning indicates the removal of the attention pruning strategy, and MLLM is not pruned; w/o Linear removes the linear compensation mechanism and replaces it with a standard residual connection; w/o Init replaces the initialization strategy of the compensation matrix $W_{c}$ proposed in Theorem 1 with zero initialization; and head layer replaces the multimodal knowledge reasoning completion layer with a conventional head-layer.
As shown in Table 3, w/o Text performance has declined significantly, while w/o Image has not declined significantly. This indicates that image tokens associated with textual context are more critical in the MKGC task. One possible explanation is that the importance of the same token may vary across different relations (as illustrated in Figure 1), a unique aspect of MKGC that $I_{image}$ representations fail to fully capture. Moreover, removing the MVTC component leads to a substantial performance drop in ELMM, primarily due to increased noise and intensified modality conflicts caused by the large number of image tokens. This ablation result further confirms that directly applying MLLMs to the MKGC task yields suboptimal outcomes. It highlights the necessity of designing targeted image token compression and selection mechanisms to mitigate cross-modal information interference specific to this task.
Notably, ELMM achieves over 30% faster inference without significant performance degradation compared to the non-pruned model, indicating redundancy in attention modules for the MKGC task. Moreover, removing the linear compensation or initializing its projection weights $W_{c}$ to zero both lead to comparable performance drops, suggesting that the effectiveness of linear compensation depends critically on appropriate initialization.
Model MR Hits@1 Hits@3 Hits@10
ELMM 112 34.1 47.3 60.7
- w/o Image 122 33.7 47.0 60.2
- w/o Text 156 30.8 43.5 57.8
- w/o MVTC 193 27.9 39.8 54.3
- w/o Pruning 106 33.8 47.2 61.0
- w/o Linear 159 33.2 44.9 58.2
- w/o Init 147 33.5 44.8 58.9
- Head Layer 121 33.9 46.9 60.6
Table 3: Ablation study results for the FB15k-237-IMG dataset. Bold and underline denote the best and second-best performance of compressed models.
4.5 Parameter Sensitivity Analysis
Figure 4: Hyperparameter analysis results. Left: Relationship between Hits@1 and the number of attention pruning layers. Right: Relationship between Hits@1 and learning rate.
4.5.1 Number of Pruning Layers
In our propose attention pruning strategy, we employ a similarity-based criterion to identify and prune the top-$K$ most similar layers. We evaluate the impact of pruning layers on performance, and the corresponding results are shown in Figure 4(left). Experimental findings indicate that the model maintains stable performance when the number of pruned layers is relatively small, implying a degree of redundancy in the attention modules of multimodal large language models (MLLMs) for MKGC. However, a marked degradation in performance is observed when the pruning depth exceeds 17 layers. Based on this analysis, we fix the pruning depth at 16 layers to achieve a favorable trade-off between inference efficiency and task accuracy.
4.5.2 Learning Rate
[…]
4.6 Impact of Backbone MLLMs on MKGC Performance
To investigate whether ELMM can continuously benefit from the rapid evolution of multimodal large language models (MLLMs), we study the impact of different backbone models on completion performance. Specifically, we select LLaVA-1.5 7B Liu et al. (2024), LLaVA-1.5 13B Liu et al. (2024), and Qwen2.5-VL 7B Bai et al. (2025a) as comparative models. The results are presented in Table 4.
From our experimental results, we observe that while LLaVA-1.5 13B contains a greater number of parameters than LLaVA-1.5 7B, the corresponding performance gains are relatively limited. This outcome may be attributed to the modest difference in parameter count between the two models. In contrast, improvements in the backbone architecture appear to have a more substantial impact on performance. For example, Qwen2.5-VL 7B consistently outperforms LLaVA-1.5 7B and even LLaVA-1.5 13B on several evaluation metrics.
Base Model MR Hits@1 Hits@3 Hits@10
LLaVA-1.5 7B 112 34.1 47.3 60.7
LLaVA-1.5 13B 108 34.8 47.7 61.0
Qwen2.5-VL 7B 87 36.7 50.1 62.2
Table 4: Results of ELMM using different base models on the FB15k-237-IMG dataset.
Model Name Inference Time MR Hits@1
ELMM 0.078 112 34.1
MKGL 0.126 172 32.5
K-ON 0.084 144 33.2
Table 5: Comparison of ELMM with state-of-the-art uni-modal methods based on large language models in terms of inference time and model performance.
4.7 Inference Time Analysis
To evaluate the inference efficiency of the proposed ELMM approach, we compare its end-to-end inference time with that of state-of-the-art uni-modal methods based on large language models, including MKGL Guo et al. (2024a) and K-ON Guo et al. (2025b), using the FB15k-237-IMG dataset. As shown in Table 5, ELMM achieves lower inference time despite the need to process multimodal information, primarily due to its attention pruning strategy. This result highlights ELMM’s advantage in inference efficiency compared to uni-modal baselines. Moreover, by effectively modeling and integrating multimodal information, ELMM also achieves superior performance across standard evaluation metrics.
5 Conclusion
We propose Efficient Lightweight Multimodal Large Language Models (ELMM), a novel framework that extends Multimodal Large Language Models (MLLMs) to the task of Multimodal Knowledge Graph Completion (MKGC). ELMM addresses two key challenges: (1) semantic noise and modality conflicts caused by excessive image tokens, and (2) the substantial computational overhead associated with processing such inputs using standard MLLMs. To mitigate these issues, we propose a Multi-view Visual Token Compressor (MVTC) based on multi-head attention mechanism, which adaptively compresses image tokens from both textual and visual views. We further apply attention pruning to remove redundant attention layers from the MLLMs backbone, improving inference efficiency. To offset any performance degradation from pruning, we add a linear projection with a carefully designed initialization scheme. Experiments on two benchmark datasets, FB15k-237-IMG and WN18-IMG, show that ELMM achieves state-of-the-art performance.